Explain how you could estimate 22% of 78 to check if your answer is reasonable.
Answer #1
22%=22/100


22% of 78= 22*78/100 = 1716/100 = 17.16
answered by: Wioleta
Answer #2
An exact answer is definitely not required.

20% is approximately a fifth (1/5)

78 is nearly 80

a fifth of 80 is 16

so the answer should be near 16
answered by: SWAPAN